OpenStudy (anonymous):

maleek bought 12 2/3 ponds of beef he used 7 1/2 ponds to make hamburgers how much ground beef does he have left maleek bought 12 2/3 ponds of beef he used 7 1/2 ponds to make hamburgers how much ground beef does he have left @Mathematics

OpenStudy (anonymous):

total - used = remaining

OpenStudy (anonymous):

so what is it

OpenStudy (anonymous):

Calculate. :-)

OpenStudy (anonymous):

can u do it 4 me

OpenStudy (anonymous):

The goal is to get you to understand to be able to do it. So, ask questions.

OpenStudy (anonymous):

if i will do this for u what u will do !!!!!!!!!! Right GT bro

OpenStudy (anonymous):

ok and i will give u a medal

OpenStudy (anonymous):

Medals are dime a dozen on this site. Learning will help you on the next problems.

OpenStudy (anonymous):

u keep ur medal i m not here to earn a medal. i love maths so m here till time i used to just give answer but after i met GT bro i have changed my mind

OpenStudy (anonymous):

You need to simply subtract two numbers. It is confusing because the numbers are in the form of mixed fractions. You need to convert each into an improper fraction first before subtracting.

OpenStudy (anonymous):

so wat is it

OpenStudy (anonymous):

How do you express 12 2/3 as a fraction?

OpenStudy (anonymous):

i dont know i just need the answer

OpenStudy (anonymous):

Do you not want to know? I can help if you do.

OpenStudy (anonymous):

i do

OpenStudy (anonymous):

12 2/3 = 12 + 2/3 = (12*3 +2)/3 = ? 7 1/2 = 7 + 1/2 = (7*2 + 1)/2 = ? Remaining = total minus used. So, you subtract the two numbers above. They are fractions. So, you need to take the common denominator and subtract.